# Q-SFT: Q-Learning for Language Models via Supervised Fine-Tuning

### v1

@@ 1 Introduction
Single-step problems, such as answering questions, can be tackled with policy gradient methods  (Ouyang et al., 2022; Rafailov et al., 2023), but sequential problems, such as dialogue or robotic control, require sample-efficient methods that can utilize data to reason about the dynamics of the problem, which typically requires training value functions (Abdulhai et al., 2023; Hong et al., 2023). Although there are many effective value-based RL methods that could be applied to LLMs and VLMs, in practice such methods have been difficult to adapt to these models with the same effectiveness as policy gradients. We posit that this is due in part to a mismatch between the pretraining objective that these models use, i.e. maximum likelihood estimation, and the fine-tuning objective necessary to train value functions. This discrepancy means that fine-tuning using multi-turn RL may require discarding some of knowledge gained by maximum likelihood pretraining of LLMs and VLMs, including a broad understanding of language, vision, and even sequential reasoning.
@@ 3 Preliminaries
Language generation can be viewed as an MDP, where states are sequences of tokens from a finite vocabulary ${\mathcal{V}}$ (Ramamurthy et al., 2023). All tokens that the agent initially observes are used as our initial state, $s_{0}=(x_{0},\ldots,x_{m})$, where $x_{i}\in{\mathcal{V}},\forall i\in[m]$. At timestep $t$, an action $a_{t}\in{\mathcal{V}}$ is some token in the vocabulary. As long as $a_{t}$ is not a special end-of-sequence <EOS> token, the transition function deterministically appends $a_{t}$ to state $s_{t}$ to form $s_{t+1}$. Otherwise, the agent observes (potentially stochastic) responses from the environment, i.e. utterances by conversational partners in the case of goal-directed dialogue, $o_{t}=(y_{0},\ldots,y_{n})$, which also consist of tokens in the vocabulary; then, the transition function appends both $a_{t}$ and responses $o_{t}$ to state $s_{t}$. This continues until the last timestep $T$ where we obtain a state $s_{T}$ and the agent receives a deterministic reward $r(s_{T})$.
@@ 4.1 Learning Values as Probabilities
Here, we see that our loss is an instance of weighted cross entropy loss with weights approximately equal to Bellman target values ^$\widehat{\mathcal{B}}^{*}p_{\bar{\theta}}(a|s)$. The primary difference is that instead of introducing a dummy action, we equally distribute the leftover weight across the remaining actions. As we will show, this acts as a label-smoothing term that ultimately regularizes the probabilities. We will show later in this section that in the absence of sampling error, our learned likelihood function $\widehat{p}_{\theta}(a|s)$ satisfies $\pi_{\beta}(a|s)\ Q^{*}(s,a)\leq\widehat{p}_{\theta}(a|s)\leq Q^{*}(s,a)$. This means that we are able to effectively learn a conservative estimation of the Q-function as a likelihood, without the need for optimizing a potentially unstable and poorly-scaling TD objective.
@@
4.2 Practical Implementation
Our objective in Equation 3 trains the model such that the predicted token probabilities match their Q-values. Our final step is to choose how to use these probabilities in the final learned policy. Prior work performs policy extraction that learns a policy regularized to be similar to the behavior policy (Peng et al., 2019; Kostrikov et al., 2021). Namely, it is well-known that a policy parameterization $\widehat{\pi}(a|s)\propto\pi_{\beta}(a|s)\exp(Q^{*}(s,a))$ is a solution to the constrained optimization problem (Peng et al., 2019; Brandfonbrener et al., 2021)
$\arg\max_{\pi}\ \mathbb{E}_{s\sim d^{\pi_{\beta}},a\sim\pi}\left[Q^{*}(s,a)\right]\quad\text{ s.t. }\quad\mathbb{E}_{s\sim d^{\pi_{\beta}}}\left[\mathrm{D}_{\mathrm{KL}}(\pi(\cdot\mid s)\ ||\ \pi_{\beta}(\cdot\mid s))\right]\leq\varepsilon.$
Recall that we have already approximated $\pi_{\beta}$ by optimizing an $\mathcal{L}_{CE}(\phi)$ over parameters $\phi$. Therefore, without any further training we can extract a policy whose probabilities of actions $\widehat{\pi}(a\mid s)$ follow $\widehat{\pi}(a|s)\propto\pi_{\phi}(a|s)\exp\left(p_{\theta}(a|s)\right)$, an expression of both our learned probabilities $p_{\theta}$ and estimated behavior policy $\pi_{\phi}$. Note that unlike prior works that explicitly require a policy extraction step with additional training, our policy can be computed at inference-time. Namely, we can express our policy using only the probabilities that we had previously learned. An overview of our entire method is provided in Algorithm 1.
Algorithm 1 Q-learning via Supervised Fine-tuning (Q-SFT)
0:  Dataset $\mathcal{D}=\{(s_{i},a_{i},r_{i},s^{\prime}_{i})\}_{i\in[N]}$
1:  Initialize $\phi,\theta,\bar{\theta}$ from pretrained model.
2:  Optimize behavior policy:
3:  for each gradient step do
4:     Update $\phi\leftarrow\phi-\alpha\nabla_{\phi}\mathcal{L}_{\mathrm{CE}}(\phi)$
5:  end for
6:  Optimize likelihood model:
7:  for each gradient step do
8:     Update $\theta\leftarrow\theta-\alpha\nabla_{\theta}\mathcal{L}_{\mathrm{WCE}}(\theta)$
10:  end for
11:  At inference time, policy probabilites become: $\widehat{\pi}(a\mid s)\propto\pi_{\phi}(a\mid s)\exp\left(p_{\theta}(a\mid s)\right)$
@@ 4.3 Theoretical Analysis
To simplify exposition, we consider a modified version of our objective in Equation 3, where instead of empirical operator $\widehat{\mathcal{B}}^{*}$, we use the true operator:

### v2

@@ 1 Introduction
Single-turn problems, such as answering questions, can be tackled with policy gradient methods  (Ouyang et al., 2022; Rafailov et al., 2023), but sequential or multi-turn problems, such as dialogue or robotic control, require sample-efficient methods that can utilize data to reason about the dynamics of the problem, which typically requires training value functions (Abdulhai et al., 2023; Hong et al., 2023). This is in multi-turn problems, the agent must plan their actions to optimize some long-term objective. Although there are many effective value-based RL methods that could be applied to LLMs and VLMs, in practice such methods have been difficult to adapt to these models with the same effectiveness as policy gradients. We posit that this is due in part to a mismatch between the pretraining objective that these models use, i.e. maximum likelihood estimation, and the fine-tuning objective necessary to train value functions. This discrepancy means that fine-tuning using multi-turn RL may require discarding some of knowledge gained by maximum likelihood pretraining of LLMs and VLMs, including a broad understanding of language, vision, and even sequential reasoning.
@@ 3 Preliminaries
Language generation can be viewed as an MDP, where states are sequences of tokens from a finite vocabulary ${\mathcal{V}}$ (Ramamurthy et al., 2023). All tokens that the agent initially observes are used as our initial state, $s_{0}=(x_{0},\ldots,x_{m})$, where $x_{i}\in{\mathcal{V}},\forall i\in[m]$. At timestep $t$, an action $a_{t}\in{\mathcal{V}}$ is some token in the vocabulary. As long as $a_{t}$ is not a special end-of-sequence <EOS> token, the transition function deterministically appends $a_{t}$ to state $s_{t}$ to form $s_{t+1}$. Otherwise, the agent observes (potentially stochastic) responses from the environment, i.e. utterances by conversational partners in the case of multi-turn dialogue, $o_{t}=(y_{0},\ldots,y_{n})$, which also consist of tokens in the vocabulary; then, the transition function appends both $a_{t}$ and responses $o_{t}$ to state $s_{t}$. This continues until the last timestep $T$ where we obtain a state $s_{T}$ and the agent receives a deterministic reward $r(s_{T})$.
@@ 4.1 Learning Values as Probabilities
Here, we see that our loss is an instance of weighted cross entropy loss with weights approximately equal to Bellman target values ^$\widehat{\mathcal{B}}^{*}p_{\bar{\theta}}(a|s)$. The primary difference is that instead of introducing a dummy action, we equally distribute the leftover weight across the remaining actions. As we will show, this acts as a label-smoothing term that ultimately regularizes the probabilities. We will show later that in the absence of sampling error, our learned likelihood function $\widehat{p}_{\theta}(a|s)$ satisfies $Q^{*}(s,a)\geq\widehat{p}_{\theta}(a|s)\geq\pi_{\beta}(a|s)\ Q^{*}(s,a)$. This means that we are able to effectively learn a conservative estimation of the Q-function as a likelihood, without the need for optimizing a potentially unstable and poorly-scaling TD objective.
@@ 4.2 Theoretical Analysis
In the previous section, we motivated a new objective $\mathcal{L}_{\mathrm{WCE}}$ given by Equation 3 that learns modified probabilities over actions $p_{\theta}$ directly from the logits of a base LLM or VLM. Here, we will show that such probabilities serve as a conservative approximation of the true Q-values. To simplify exposition, we consider a simple modification of Equation 3, where instead of empirical operator $\widehat{\mathcal{B}}^{*}$, we use the true operator:
@@
4.3 Practical Implementation
Our objective in Equation 3 trains the model such that the predicted token probabilities match their Q-values. Our final step is to choose how to use these probabilities in the final learned policy. Prior work performs policy extraction that learns a policy regularized to be similar to the behavior policy (Peng et al., 2019; Kostrikov et al., 2021). Namely, it is well-known that a policy parameterization $\widehat{\pi}(a|s)\propto\pi_{\beta}(a|s)\exp(\beta\,Q^{*}(s,a))$ is a solution to the constrained optimization problem (Peng et al., 2019; Brandfonbrener et al., 2021)
$\arg\max_{\pi}\ \mathbb{E}_{s\sim d^{\pi_{\beta}},a\sim\pi}\left[Q^{*}(s,a)\right]\quad\text{ s.t. }\quad\mathbb{E}_{s\sim d^{\pi_{\beta}}}\left[\mathrm{D}_{\mathrm{KL}}(\pi(\cdot\mid s)\ ||\ \pi_{\beta}(\cdot\mid s))\right]\leq\varepsilon\,,$
where $\beta>0$ is a hyperparameter derived from the Lagrange multiplier. Recall that we have already approximated $\pi_{\beta}$ by optimizing an $\mathcal{L}_{CE}(\phi)$ over parameters $\phi$. Therefore, without any further training we can extract a policy whose probabilities of actions follow $\widehat{\pi}(a|s)\propto\pi_{\phi}(a|s)\exp\left(\beta\,p_{\theta}(a|s)\right)$, which can be computed at inference time using our learned probabilities $p_{\theta}$, estimated behavior policy $\pi_{\phi}$, and tunable hyperparameter $\beta>0$. Note that unlike prior works that explicitly require a policy extraction step with additional training, our policy can be computed at inference-time. Namely, we can express our policy using only the probabilities that we had previously learned. An overview of our entire method is provided in Algorithm 1. We also provide implementation details such as hyperparameter selection for our experiments in Appendix B.
Algorithm 1 Q-learning via Supervised Fine-tuning (Q-SFT)
0:  Dataset $\mathcal{D}=\{(s_{i},a_{i},r_{i},s^{\prime}_{i})\}_{i\in[N]}$, hyperparameter $\beta>0$
1:  Initialize $\phi,\theta,\bar{\theta}$ from pretrained model.
2:  Optimize behavior policy:
3:  for each gradient step do
4:     Update $\phi\leftarrow\phi-\lambda_{\phi}\nabla_{\phi}\mathcal{L}_{\mathrm{CE}}(\phi)$
5:  end for
6:  Optimize likelihood model:
7:  for each gradient step do
8:     Update $\theta\leftarrow\theta-\lambda_{\theta}\nabla_{\theta}\mathcal{L}_{\mathrm{WCE}}(\theta)$
10:  end for
11:  At inference time, policy probabilites become: $\widehat{\pi}(a\mid s)\propto\pi_{\phi}(a\mid s)\exp\left(\beta\,p_{\theta}(a\mid s)\right)$
@@ Appendix A Theoretical Proofs
Appendix B Implementation Details
We use the hyperparameters reported in Table 4. All algorithms were trained on a single TPUv3 on Google Cloud until convergence.
Hyperparameter Chess Wordle 20Q WebShop ALFWorld
$\beta$ 8.0 4.0 1.0 1.0 1.0
$\gamma$ discount factor 0.99 0.99 0.95 0.9 0.95
Batch size 128 128 128 128 128
Target network update $\alpha$ 0.005 0.005 0.005 0.01 0.01
Number of updates per iteration 60 60 60 50 50
Number of iterations 100 100 100 200 200
$\lambda_{\phi}$ learning rate 1e-4 1e-4 1e-4 2e-4 3e-4
$\lambda_{\theta}$ learning rate 1e-4 1e-4 1e-4 2e-4 1e-4
Table 4: Hyperparameters used during training Q-SFT in our experiments.
As shown in Table 4, most hyperparameters were held the same except for $\beta$, where larger $\beta$ results in a more deterministic policy. In practice, we only had to increase $\beta$ for tasks with restricted action spaces (such as games). Hence, we can conclude that in most practical tasks, our method does does not require much hyperparamter tuning to perform well.